OpenStudy (richyw):

precalc question. How to I work back from a fraction to find its angle? I have like \[\frac{1+\sqrt{3}}{2\sqrt{2}}\]\[\frac{-1+\sqrt{3}}{2\sqrt{2}}\]How am I supposed to see this and figure out that they are \(\cos(\pi /12)\text{ and } \sin(\pi /12)\) respectively?

OpenStudy (precal):

using a half angle, perhaps

OpenStudy (precal):

you know cosine 30 degrees and sine 30 degrees, I would use a half angle identity to find this value

OpenStudy (precal):

I could be way off base, let's see what electrokid has to say

OpenStudy (anonymous):

express the given fraction interms of known t-ratio values \[{1+\sqrt{3}\over 2\sqrt{2}}={1\over\sqrt{2}}\cdot{1\over2}+{\sqrt{3}\over2}\cdot{1\over\sqrt{2}}\\ \quad=\sin(\pi/4)\cdot\cos(\pi/3)+\sin(\pi/3)\cos(\pi/4)\\ \quad=\sin(\pi/4+\pi/3) \]

OpenStudy (anonymous):

@precal thanq :)

OpenStudy (precal):

this is the sum and difference formulas